OpenStudy (anonymous):

Tristan walks 3over4 of a mile to school every day. He wants to find how long it will take him to walk 6 miles. Use the model to find 6 ÷ 3over4, the number of days it will take Tristan to walk 6 miles.

OpenStudy (bahrom7893):

So Tristan walks: 3/4 mi - 1 day 6 mi - ? days Cross multiply: ? = 6*1 / (3/4)

OpenStudy (anonymous):

so whats the answer ?

OpenStudy (bahrom7893):

6/3/4 = 6*4/3 = 24/3 = 8

OpenStudy (anonymous):

so its 8

OpenStudy (anonymous):

hello?